OpenStudy (anonymous):

The equation of the line with x-intercept −2 and y-intercept 5 can be written in the form y=mx+b what are m and b?

OpenStudy (anonymous):

m is the gradient/slop b is the y-intercept

OpenStudy (anonymous):

*slope

OpenStudy (anonymous):

what are the figures for m and b in this case?

OpenStudy (anonymous):

well, b is 5 (says in the equation) now you need to find m

OpenStudy (anonymous):

the gradient formula is: m=(y2-y1)/(x2-x1)

OpenStudy (anonymous):

but i dont have points. i just have x and y intercepts

OpenStudy (anonymous):

they are points x intercept: (-2,0) y intercept: (0,5)

OpenStudy (anonymous):

you need to realise that when it says x-intercept, it means that y is zero, and vice versa. This is because when you draw it on a plane. You'll find the x-intercept is where y=0 and x is on the axis

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

one sec let me attempt

OpenStudy (anonymous):

5/2

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thanks can you help me with this last question please

OpenStudy (anonymous):

Find an equation y=mx+b of the perpendicular bisector of the line segment joining the points A(7,7) and B(13,1). i just need to find the constant b

OpenStudy (anonymous):

I do not understand the whole thing, I mean I think I do but I can't be sure. I'll tell you what I know though. to find b, you need to use the gradient and one of the points (any one will do)

OpenStudy (anonymous):

so using this equation: \[y-y_1=m(x-x_1)\] you can find out what b is by subing in the values of m, and one of the coordinate points

OpenStudy (anonymous):

wait, I think that's wrong. Ignore that

OpenStudy (anonymous):

i got it nvm

OpenStudy (anonymous):

so, here's what i think you have to do. Find the midpoint of that line segment using: \[M=(\frac{ x_1+x_2 }{ 2 },\frac{ y_1+y_2 }{ 2 })\] then find the gradient of the line with: \[m=\frac{ y_2-y_1 }{ x_2-x_1 }\] then using that coordinate point you just found use: \[y-y_1=m(x-x_1)\] to find the equation of the line